OpenStudy (anonymous):

Q- 7 5/6 = 6 1/2 is -1/4?

hero (hero):

\[Q - 7 \frac{5}{6} = 6 \frac{1}{2}\] \[Q = 6 \frac{1}{2} + 7 \frac{5}{6}\]

OpenStudy (anonymous):

I am confused on how I would solve this

OpenStudy (anonymous):

Its the top 1

hero (hero):

@ellie67, the "bottom one" is what you get when you add seven and five-sixths to both sides. In order to solve this, you have to isolate Q.

hero (hero):

Which means, in order to solve this, you have to add both fractions together.

OpenStudy (anonymous):

so isolate Q add both factions would that make it 121/2?

hero (hero):

If you add correctly you get: \[6 + \frac{3}{6} + 7 + \frac{5}{6}\] \[=(6 + 7) + \left(\frac{3}{6} + \frac{5}{6}\right)\] \[=13 + \frac{8}{6}\] \[=13 + \frac{6}{6} + \frac{2}{6}\] \[=13 + 1 + \frac{1}{3}\] \[=14 + \frac{1}{3}\] \[=14 \frac{1}{3}\]